OpenStudy (anonymous):

Having trouble with my homework the question is find x in : x+1 / 5 = 2x- 2/ 6 PLEASE HELP

OpenStudy (anonymous):

x would be equal to 4

OpenStudy (anonymous):

Thank you did you divide by six first or what did you do first)

OpenStudy (anonymous):

take 2 common from the numerator of left hand side and divide it with 6 to simplify

OpenStudy (anonymous):

what are the steps when you have two fractions

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

results in the new equation x+1/5 =x-1/3

OpenStudy (anonymous):

now cross multiply the equations (whenever you have two fractions)

OpenStudy (anonymous):

makes a new equation 3(x+1)= 5(x-1) by cross multiplying

OpenStudy (anonymous):

now simply multiply the 3 and 5 with their respective contents in the brackets. end result x=4

OpenStudy (anonymous):

hope this gives clear idea to you

OpenStudy (anonymous):

YES !! thank you :) it's surprisingly clear now :) have a good night.

OpenStudy (anonymous):

no problem! :)

OpenStudy (anonymous):

one more thing just to clarify the 2 common are the common factors ? or what ? I knew like five seconds ago...